OpenStudy (anonymous):

1. (2x2 + y)(2x2 - y) 2. 4(2x2 + y)(2x2 - y) 3. (x + 4y)(x - 4y) 4. 4(x + 4y)(x - 4y) 5. prime a. Factor x2 - 16y2 b. Factor 4x4 - y2 c. Factor 4x2 - 64y2 d. Factor 9x2 - 8y2 e. Factor 16x4 - 4y2 Can someone please help me with this?!?!?!

OpenStudy (anonymous):

What's number 5?

OpenStudy (anonymous):

I can do the first two for you of each 1-5 and a and b. You can figure out the rest from there, hopefully.

OpenStudy (anonymous):

1. (2x2 + y)(2x2 - y) This is a called a "difference of squares", something in the form of (ax2 + x)(ax2 - x) can simply be written as: \[\left( ax - x \right)^{2}\] With that being said, we can say that the first problem is simply: \[\left( 2x - y \right)^{2}\] 2. Same thing, just leave the 4 outside: \[4* \left( 2x-y \right)^{2}\] Hopefully you can see the pattern. a) x2 - 16y2. (x-4y)(x+4y) b) 4x4 - y^2 (2x2 - y)(2x2 + y)